OpenStudy (volleyballlover55):

A standard showerhead in Andrew's house dispenses 8 gallons of water per minute. Andrew changed this showerhead to an energy-saving one. The equation shows the amount of water dispensed, y, as a function of the number of minutes, x, for the new showerhead. y = 3x How much water does Andrew save each day with the change in showerheads if he uses the shower for 8 minutes each day? 5 gallons 16 gallons 24 gallons 40 gallons

OpenStudy (volleyballlover55):

@Michele_Laino

OpenStudy (michele_laino):

the requested quantity, is given by the subsequent difference: amount of water with \(old\) showerhead - amount of water with the \(new\) showerhead= = \((8 \times 8) - (3 \times 8)...?\)

OpenStudy (volleyballlover55):

64-24 = 40

OpenStudy (michele_laino):

correct!

OpenStudy (volleyballlover55):

Ada and Steve spend a certain amount of money from their accounts each week at a pet shelter. The table shows the relationship between the amount of money (y) remaining in Ada's account and the number of weeks (x): Function 1: Number of Weeks (x) Amount Remaining (dollars) (y) 1 40 2 36 3 32 4 28 The equation shows the relationship between the amount of money, y, remaining in Steve's account and the number of weeks, x: Function 2: y = –5x + 40 Which statement explains which function shows a greater rate of change? Function 1, because Ada spends $4 each week and Steve spends –$5 each week Function 1, because Ada spends $12 each week and Steve spends $35 each week Function 2, because Steve spends $5 each week and Ada spends $4 each week Function 2, because Steve spends $40 each week and Ada spends $12 each week